### v1

@@ 3.10 Experimental Results and Performance Validation
Cross-Model Universality and Architecture-Agnostic Validation: Our comprehensive evaluation across five diverse transformer architectures—ranging from compact GPT-2 (125M) to larger language models (LFM2 at 2.6B parameters)—demonstrates that QEIL’s scaling laws and heterogeneous orchestration principles hold universally, independent of specific model architecture. All models achieve 66.5%–70.0% pass@k coverage with energy-aware execution compared to 56–63% baseline, representing 7–10.5 percentage point improvements consistent across the model landscape. This universality validates our core hypothesis: inference-time scaling laws are architecture-agnostic, parameterized only by model size $N$, sample budget $S$, tokens per sample $T$, and precision $Q$. Energy improvements remain substantial across diverse scales—from 47.7% reduction for GPT-2 to 35.6%–78.2% reduction for larger models—confirming that heterogeneous orchestration benefits scale across model families.
@@
Coverage Improvement and Sample Efficiency (Figure 4): Our energy-aware orchestration achieves 66.5%–70.0% pass@k coverage across all models with 20 samples, compared to 56%–63% baseline coverage with standard homogeneous inference—representing 7–10.5 percentage point improvements. This demonstrates that heterogeneous task-device matching enables superior sample aggregation by reducing per-sample latency variance, optimizing computation paths for each inference stage, and leveraging device-specific arithmetic intensity characteristics (prefill on GPU for high-compute phases, decode on NPU for bandwidth-bound phases). The coverage vs. sample count analysis validates Theorem 1.1 (coverage scaling law), confirming that intelligent device placement allows more effective utilization of each sample generation. The consistency across models—all achieving 70% energy-aware coverage except Qwen2 at 66.5%—demonstrates that the scaling law exponent $\beta\approx 0.7$ holds universally across architectures.
Multi-Sample Aggregation Efficiency (Figure 3): The multi-sample aggregation efficiency reveals a critical insight: heterogeneous orchestration decouples per-sample latency from aggregate coverage by enabling efficient sample parallelization and device-specific optimization. Energy-aware configurations achieve 66.5%–70.0% pass@k coverage across all models, with improvements ranging from +7.0 percentage points for Llama-3.2-1B to +10.5 percentage points for GPT-2 and Qwen2-0.5B. Models with lower baseline coverage (Qwen2 at 56%, GPT-2 at 59.5%) achieve larger absolute improvements, consistent with logarithmic scaling where initial samples provide highest marginal value. The inverse relationship between baseline coverage and improvement magnitude demonstrates that heterogeneous orchestration disproportionately benefits models with lower initial performance, enabling “catch-up” on weaker architectures through intelligent device routing that maximizes information content per sample.
Latency and Throughput Improvements: Beyond energy and coverage, heterogeneous orchestration simultaneously improves latency and throughput. End-to-end inference latency improves by average 15.8% (ranging from 8.0%–22.5% across models), with specific improvements: 22.5% for GPT-2 (1.73 ms to 1.34 ms), 16.6% for Granite-350M (1.69 ms to 1.41 ms), 8.0% for Qwen2-0.5B (1.76 ms to 1.62 ms), 13.1% for Llama-3.2-1B (1.91 ms to 1.66 ms), and 18.8% for LFM2-2.6B (1.86 ms to 1.51 ms). This latency improvement stems from optimized memory bandwidth utilization on device-specific hardware: NPUs and GPUs with 200–900 GB/s bandwidth dramatically outpace CPUs (30–100 GB/s) for memory-bound decode operations, reducing per-sample latency. The latency gains are inversely correlated with model size—smaller models (GPT-2, Granite) achieve larger latency reductions due to better NPU utilization, while larger models (LFM2) still achieve substantial gains through GPU-NPU pipelining.
Price-Power-Performance (PPP) Trade-off and Economic Viability: The PPP score—capturing cost-power-performance balance—improves by average 39.0% across models, with energy-aware configurations achieving 20.74 PPP for GPT-2 (versus 16.85 baseline, +23.1%), 16.60 PPP for Granite-350M (versus 10.90 baseline, +52.3%), and 15.49 PPP for Qwen2-0.5B (versus 10.83 baseline, +43.0%). This metric reflects the composite economic advantage: heterogeneous deployment amortizes hardware costs (by spreading computation across cheaper NPU and CPU resources), reduces electricity costs (through 35%–78% energy savings), and maintains throughput. The PPP metric is particularly valuable for evaluating practical deployment viability at scale, where even marginal cost-per-query reductions accumulate to substantial savings across millions of inferences. PPP improvements of 23%–52% demonstrate that heterogeneous orchestration is economically superior to cloud-centric inference for edge deployment, with LFM2-2.6B achieving exceptional PPP gains (+32.8%) despite minimal IPW improvement.
Power Consumption Reduction and Thermal Viability: Average power consumption drops dramatically across all models: 83.5 W for GPT-2 (from 402.5 W baseline, $79.2\%$), 82.3 W for Granite-350M (from 460.4 W, $82.1\%$), 74.4 W for Qwen2-0.5B (from 244.7 W, $69.6\%$), 79.0 W for Llama-3.2-1B (from 164.5 W, $52.0\%$), and 75.0 W for LFM2-2.6B (from 175.8 W, $57.3\%$). This mean power reduction of 68.0% is transformative for edge deployment: heterogeneous orchestration reduces average power consumption to the 75–84 W range, fitting within stringent thermal budgets typical of edge devices (65–125 W TDP for passively-cooled or fan-less hardware). This thermal reduction eliminates cooling infrastructure requirements—a major cost and reliability factor in edge systems—enabling deployment of large language model inference on consumer-grade IoT devices, mobile edge servers, and battery-powered applications where datacenter-scale GPUs are thermally infeasible. The power reduction is most dramatic for models originally demanding high compute (Granite-350M, GPT-2 with 402–460 W baselines), indicating that heterogeneous orchestration disproportionately benefits compute-heavy inference patterns.
Energy Consumption Analysis and Practical Constraints: Total energy consumption for 20 inference samples shows substantial reductions consistent across models: GPT-2 achieves 47.7% energy reduction (43.1 kJ to 22.5 kJ), Granite-350M achieves 78.2% reduction (403.1 kJ to 88.0 kJ), Qwen2-0.5B achieves 46.7% reduction (352.3 kJ to 187.9 kJ), Llama-3.2-1B achieves 35.6% reduction (330.5 kJ to 213.0 kJ), and LFM2-2.6B achieves 35.9% reduction (490.3 kJ to 314.3 kJ). The mean energy reduction of 48.8% directly translates to extended battery lifetime for edge devices: a device with 1 MJ energy budget can execute approximately 44 energy-aware inferences versus 23 standard inferences for GPT-2, doubling query capacity. Energy improvements scale inversely with model size for some metrics (Granite-350M achieves 78.2%, GPT-2 achieves 47.7%), suggesting that mid-range models benefit most from heterogeneous optimization, while very large models (LFM2-2.6B) achieve more modest but still substantial energy gains through different optimization mechanisms.
@@
Figure 4: Coverage scaling and sample efficiency analysis across five model families showing consistent 66.5%–70.0% pass@k coverage achieved by energy-aware orchestration at $N=20$ samples, compared to 56%–63% baseline coverage across models, demonstrating 7–10.5 percentage point improvements through heterogeneous orchestration. The universality of the coverage improvement across diverse architectures (GPT-2, Granite, Qwen, Llama, LFM) validates Theorem 1.1—that coverage scaling law exponent $\beta\approx 0.7$ is architecture-agnostic and holds universally across all transformer families.
4 Conclusion
This paper presents QEIL, a comprehensive framework for quantifying and optimizing edge inference through principled scaling law characterization combined with heterogeneous device orchestration. By decomposing the inference pipeline into device-aware tasks and leveraging five fundamental theorems governing coverage, energy, latency, cost, and device-task efficiency, we demonstrate that intelligent hardware-layer mapping achieves transformative improvements over homogeneous inference approaches across diverse model families and architectures.
[…]
The experimental results on our heterogeneous edge platform (Intel Core Ultra 9 285HX processor with integrated Intel AI Boost NPU, NVIDIA RTX PRO 5000 Blackwell GPU, and Intel Graphics GPU) validate that heterogeneous orchestration fundamentally outperforms cloud-based homogeneous inference for latency-sensitive and power-constrained edge deployments. Notably, the universality of improvements across five distinct model architectures—GPT-2, Granite, Qwen, Llama, and LFM families—demonstrates that inference-time scaling laws are architecture-agnostic and that heterogeneous optimization principles generalize across the model landscape from 125M to 2.6B parameters. By simultaneously improving pass@k coverage, reducing total energy consumption by up to 78.2%, cutting average power consumption by 68% (to 75–84 W range suitable for edge thermal budgets), and improving latency by 15.8% on average, QEIL enables practical edge intelligence within strict device resource budgets while maintaining competitive accuracy-efficiency trade-offs.
@@ 4 Conclusion
Future work would benefit from evaluation on more diverse heterogeneous infrastructure including additional NPU architectures (Qualcomm Snapdragon NPU, MediaTek Dimensity NPU), mobile edge platforms (NVIDIA Jetson Orin, Google Tensor Processing Units), and emerging ASIC-based accelerators (GraphCore IPUs, Groq LPUs). Extending QEIL to support dynamic reallocation strategies responding to runtime device state changes (thermal throttling, power anomalies, memory pressure) would further enhance the framework’s robustness in production environments. Additionally, investigating distributed inference across multiple edge nodes and quantifying inter-device communication overhead would address deployment scenarios in edge computing clusters, particularly for bandwidth-constrained IoT networks and mobile networks with variable connectivity. Exploring integration with model compression techniques (quantization, pruning, distillation) and alternative LLM architectures (state-space models, mixture-of-experts) would expand QEIL’s applicability to emerging model families and further optimize energy consumption.
@@
QEIL represents a significant step toward principled, measurement-driven approaches to edge inference optimization, providing both theoretical foundations (architecture-agnostic scaling law theorems) and practical tools (greedy layer assignment, dynamic orchestration, unified efficiency metrics) for system designers to deploy large language models on resource-constrained devices while achieving competitive accuracy-efficiency trade-offs. The demonstrated universality across five model families, consistent improvements across diverse architectures (7–10.5 percentage points coverage gain, 48.8% mean energy reduction, 68% mean power reduction), and practical economic viability (39% PPP improvement) establish heterogeneous edge computing as the optimal deployment strategy for energy-constrained AI systems, shifting the paradigm from centralized cloud inference to intelligent local edge execution.

### v2

@@ 5.1 Inference-Time Scaling Behavior Across Heterogeneous Hardware
Beyond coverage, we analyze the scaling of energy consumption, latency, and cost. We find that these quantities exhibit distinct but systematic scaling trends, reflecting the interaction between model complexity and device-specific hardware constraints. These observations indicate that inference-time efficiency is shaped not only by model parameters but also by the underlying hardware execution regime.
5.2 Impact of Heterogeneous Orchestration on Inference Efficiency
We next evaluate the effect of heterogeneous task allocation across CPUs, GPUs, and NPUs. Across all evaluated models, heterogeneous orchestration consistently outperforms homogeneous execution on the best single device, yielding simultaneous improvements in energy efficiency, latency, and coverage.
Quantitatively, heterogeneous execution increases Intelligence Per Watt by 4.8–5.6× while reducing total energy consumption by 47–78% relative to homogeneous baselines. Importantly, these gains are achieved without sacrificing model accuracy. This suggests that hardware heterogeneity is not merely an implementation detail but a fundamental factor governing inference-time efficiency.
5.3 Multi-Objective Trade-offs Between Coverage, Energy, and Latency
Inference-time optimization inherently involves trade-offs between coverage, energy, and latency. To systematically analyze these trade-offs, we introduce composite efficiency metrics—Intelligence Per Watt (IPW), Energy-Coverage Efficiency (ECE), and Price-Power-Performance (PPP). These metrics enable principled comparison across heterogeneous configurations.
Across models, QEIL consistently shifts the Pareto frontier of inference performance, achieving higher coverage at lower energy and latency than homogeneous baselines. This indicates that heterogeneous orchestration enables a qualitatively different optimization regime, rather than incremental improvements within an existing regime.
5.4 Robustness Across Model Families and Hardware Configurations
To assess generality, we evaluate QEIL across multiple model families and hardware configurations. We observe that the qualitative scaling behavior remains consistent across models spanning two orders of magnitude in parameter count. Moreover, the relative benefits of heterogeneous orchestration persist across diverse device combinations, including CPU–GPU and GPU–NPU configurations.
These results suggest that the observed scaling relationships are not specific to a single model or device but reflect broader structural properties of inference-time computation in transformer-based systems.
5.5 Edge vs Cloud Inference Regimes
Finally, we analyze the relative efficiency of edge and cloud inference. We observe a transition between regimes in which heterogeneous edge execution becomes more energy-efficient for small-to-medium model scales, while homogeneous cloud execution dominates at larger scales. This transition highlights the existence of distinct inference regimes governed by model scale and hardware constraints.
Figure 6: Coverage scaling and sample efficiency analysis across five model families showing consistent 66.5%–70.0% pass@k coverage achieved by energy-aware orchestration at $N=20$ samples, compared to 56%–63% baseline coverage across models, demonstrating 7–10.5 percentage point improvements through heterogeneous orchestration. The consistency of the coverage improvement across diverse architectures (GPT-2, Granite, Qwen, Llama, LFM) validates Scaling Formalism 1.1—that coverage scaling exponent $\beta\approx 0.7$ generalizes across all tested transformer families.
Collectively, these findings suggest that inference-time scaling and heterogeneous orchestration jointly define a new design space for efficient AI deployment.
@@ 5.5 Edge vs Cloud Inference Regimes
6 Conclusion
This paper presents QEIL, a comprehensive framework for quantifying and optimizing edge inference through principled scaling characterization combined with heterogeneous device orchestration and safety-first design. By decomposing the inference pipeline into device-aware tasks and leveraging five empirically validated scaling formalisms governing coverage, energy, latency, cost, and device-task efficiency, we find that intelligent hardware-layer mapping achieves consistent improvements over homogeneous inference approaches across diverse model families and architectures.
@@ 6 Conclusion
The experimental results on our heterogeneous edge platform (Intel Core Ultra 9 285HX processor with integrated Intel AI Boost NPU, NVIDIA RTX PRO 5000 Blackwell GPU, and Intel Graphics GPU) suggest that heterogeneous orchestration consistently outperforms cloud-based homogeneous inference for latency-sensitive and power-constrained edge deployments. Notably, the consistency of improvements across five distinct model architectures—GPT-2, Granite, Qwen, Llama, and LFM families—indicates that inference-time scaling formalisms generalize across the transformer model landscape from 125M to 2.6B parameters. By simultaneously improving pass@k coverage, reducing total energy consumption by up to 78.2%, cutting average power consumption by 68% (to 75–84 W range suitable for edge thermal budgets), improving latency by 15.8% on average, and maintaining safe thermal operation throughout, QEIL enables practical edge intelligence within strict device resource budgets while maintaining competitive accuracy-efficiency trade-offs.
@@
Building on the foundational insights of Asgar et al. Asgar et al. (2025) regarding heterogeneous AI infrastructure for datacenter-scale agentic workloads, QEIL extends these formalisms to resource-constrained edge computing environments. Our work demonstrates that agentic orchestration across multi-vendor heterogeneous hardware—whether from the same vendor (Intel CPU + Intel NPU + Intel GPU) or different vendors (Intel processors + NVIDIA GPU)—enables practical edge AI deployment with simultaneous improvements in efficiency, coverage, latency, and safety. The cross-dataset validation on WikiText-103, GSM8K, and ARC-Challenge confirms that our scaling formalisms are task-agnostic within the transformer family.
Critically, our safety-first design philosophy demonstrates that reliability and efficiency are complementary rather than competing objectives. By implementing thermal protection that proactively manages device temperatures, fault tolerance that enables graceful degradation under hardware failures, and adversarial robustness that prevents malicious inputs from destabilizing the system, QEIL bridges the gap between research prototypes and production-ready edge AI systems. The 3-5% efficiency overhead of safety mechanisms is more than offset by the elimination of thermal throttling events and the assurance of predictable, reliable operation—essential requirements for consumer-facing applications where device damage or system crashes destroy user trust.
@@
Future work would benefit from evaluation on more diverse heterogeneous infrastructure including additional NPU architectures (Qualcomm Snapdragon NPU, MediaTek Dimensity NPU), mobile edge platforms (NVIDIA Jetson Orin, Google Tensor Processing Units), and emerging ASIC-based accelerators (GraphCore IPUs, Groq LPUs). Extending QEIL to support dynamic reallocation strategies responding to runtime device state changes (thermal throttling, power anomalies, memory pressure) would further enhance the framework’s robustness in production environments. Additionally, investigating distributed inference across multiple edge nodes and quantifying inter-device communication overhead would address deployment scenarios in edge computing clusters, particularly for bandwidth-constrained IoT networks and mobile networks with variable connectivity. Exploring integration with model compression techniques (quantization, pruning, distillation) would expand QEIL’s applicability to emerging model families and further optimize energy consumption. Testing on non-transformer architectures (CNNs, diffusion models, graph neural networks) would help establish whether scaling exponents generalize beyond the transformer family. Finally, extending the safety framework to cover additional failure modes (memory corruption, driver crashes, adversarial model attacks) would strengthen QEIL’s suitability for safety-critical applications.
@@
QEIL represents a step toward principled, measurement-driven approaches to edge inference optimization, providing both empirically grounded scaling formalisms and practical tools (greedy layer assignment, dynamic orchestration, unified efficiency metrics, safety monitoring) for system designers to deploy large language models on resource-constrained devices while achieving competitive accuracy-efficiency trade-offs. The demonstrated generalizability across five model families, consistent improvements across diverse architectures (7–10.5 percentage points coverage gain, 48.8% mean energy reduction, 68% mean power reduction), practical economic viability (39% PPP improvement), and robust safety guarantees (zero thermal events, 100% fault recovery) establish heterogeneous edge computing as an effective deployment strategy for energy-constrained AI systems, complementing centralized cloud inference with intelligent local edge execution. By treating the heterogeneous orchestrator as an intelligent agentic optimizer responsible for both performance optimization and system health, QEIL demonstrates that “safety-first, capability-second” design enables rather than constrains practical edge AI deployment.